OpenStudy (anonymous):

e^x/5 - e^-x/5 how do you factor this out? i believe you take out e^-x/5, then...what next

OpenStudy (anonymous):

e^-x = 1/e^x

OpenStudy (anonymous):

Now take LCM of the denominators and go ahead

OpenStudy (anonymous):

oh ok thanks :)

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

wait...u take out e^-x?

OpenStudy (anonymous):

not e^-x/5?

OpenStudy (anonymous):

OK. \[\large \frac{e^x}{5} - \frac{1}{5e^x}= \frac{1}{5e^x}{(e^{2x} - 1)}\]

OpenStudy (anonymous):

srry..it is e^ (x/5) ..forgot my brackets

OpenStudy (anonymous):

e^ (x/5) - e (-x/5)

OpenStudy (anonymous):

Apply same logic :D

OpenStudy (anonymous):

i took out e^(-x/5)

OpenStudy (anonymous):

keep typing it wrong..e^ (x/5) - e ^(-x/5)

OpenStudy (anonymous):

so i got e^ (-x/5) ( e^-1 - 1) is this right?

OpenStudy (anonymous):

Not exactly \[\large {e^{x/5}} - \frac{1}{e^{x/5}}= \frac{1}{e^{x/5}}{(e^{2x/5} - 1)}\]

OpenStudy (anonymous):

oh okok i get it thanks

OpenStudy (anonymous):

Welcome :D